OpenStudy (anonymous):

Hey this is probably easy I know it is but this is just a checkup to see if Im right.. I think the answer is 32 because if x=4 and y=8 then 8x4=32! right or wrong?? here is the question and answer incase you need it!... There is an inverse relationship between x and y. If y is sixteen when x is 4, what is x when y is 8? 2 8 16 32

OpenStudy (anonymous):

Ummm... It's actually 8. Inverse means x*y is constant. So 8x=16*4. Therefore, 8x=64 so x=8.

OpenStudy (anonymous):

@vipergirl81 omg I feel so stupid... thanks for correcting me.

OpenStudy (anonymous):

X=8

OpenStudy (jdoe0001):

or worded differently, the INVERSE RELATION will be \(\bf y = \cfrac{n}{x}\) what is the "constant of variation"? x * y = n